Jamie has 6 quarters and some dimes in his pocket. the total value of the coins is 4.50. How many dimes does he have in his pocket?
Answer by TimothyLamb(4379)

x = number of dimes
---
6*25 + 10x = 450
150 + 10x = 450
10x = 450 - 150
x = (450 - 150)/10
x = 30 dimes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php